Three equal charges of -3.2x10^-6 C are situated at the corners of a triangle of side 3 m. Find the electric field at the center of the triangle.


1
Expert's answer
2021-05-12T16:43:32-0400

Answer

Magnitude of electric field due to each charge will be equal so

"E=\frac{(9\times10^9)(3.2\times10^{-6})}{(\frac{3}{\sqrt{2}})^2}"


"E=6.4\times10^3N/C"



Net electric field at center

"E_{net}=6.4\times10^3(2cos30°-1) \\E_{net}=4.68\times10^{3}N/C"